fill the empty spaces
f Cumulative % Cumulative% 50.0 129 Language Proficiency 1 Language 2 Languages 3+ Languages I 214 T
Answer #1

f 129 Language Proficiency 1 Language 2 Language 3+ Language Cumulative f 129 214 T 258 % 50.00 32.95 17.05 Cumulative % 50.0

First we calculate total using cumulative class %.

class % for 1st class = cumulative % for first class = 50% and

class % for 1st class = class f for 1st class / Total = 50%

or, class f for 1st class/Total = 0.5

or, Total = class f for 1st class/ 0.5 = 129 / 0.5 = 258

Now, we get the class frequency for 2nd class.

class f for 2nd class = cumulative f for 2nd class - class f for 1st class

or, class f for 2nd class = 214- 129 = 85

Now, we get the class frequency for 3rd class.

class f for 3rd class = Total - class f for 2nd class - class f for 1st class

class f for 3rd class = 258 - 85 - 129 = 44.

Now we calculate class % for each class:

For 1st class we have already calculated it.

class % for 2nd class = class f for 2nd class / total

or, class % for 2nd class = 85 / 258 = 32.95

Similarly class % for 3rd class = 44 / 258 = 17.05

Cumulative f is calculated as:

Cumulative f for any class = sum of class f upto that class

so,

Cumulative f for 1st class = 129

Cumulative f for 2nd class = 129 + 85 = 214

Cumulative f for 3rd class = 129 + 85 + 44 = 258

Cumulative % for any class = sum of class % upto that class

so,

Cumulative % for 1st class = 50

Cumulative % for 2nd class = 50 + 32.95 = 82.95

Cumulative % for 3rd class = 50 + 32.95 + 17.05 = 100